OpenStudy (anonymous):

For what values of a is sqrt{10-a} defined?

myininaya (myininaya):

can the inside be negative?

myininaya (myininaya):

can we have sqrt{negative number}?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\sqrt{a-10}\]is defined if the radicand is greater than or equal to zero

myininaya (myininaya):

so we want the inside to be positive or zero which means we want it to be > or = 0

myininaya (myininaya):

\[10-a \ge 0\]

OpenStudy (anonymous):

can we have \[\sqrt{\text{negative number}}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no